OpenStudy (anonymous):

Amplitude 2, period pi/4, poiny (1/6,0)

OpenStudy (anonymous):

point

OpenStudy (anonymous):

what are you exactly looking for?

OpenStudy (ranga):

Find the function? Are they asking for sine or the cosine function?

OpenStudy (anonymous):

Write an equation for a sine curve that has the given amplitude and period, and which passes through the given point.

OpenStudy (jdoe0001):

so... what do you think, anything you can pad sin(x) with the provided information? just whatever you can pad it, doesn't have to be much

OpenStudy (ranga):

Okay. The general sine function is y = Asin(Bx + C) where A is the Amplitude and 2(pi)/B is the period. y = 2sin(Bx + C) and 2(pi)/B = (pi)/4. So B = 8 y = 2sin(8x + C) It has the point (1/6, 0). Put x = 1/6 and y = 0 and find C.

OpenStudy (ranga):

The period for the function shown in the link is (pi)/2 not (pi)/4 as stated in this problem.